# Energy-dependent intrinsic time delay of gamma-ray bursts on testing Lorentz invariance violation

### v1

@@ 3.1 Model A
We use the same GRBs data as [21, 22]. As shown in Fig. 2 of [22], the distribution of those high-energy photons seems to show a regularity that they can be considered in two cases: the mainline photons case (9 high-energy photons are involved) and the all photons case (all 14 high-energy photons are involved). We also consider the same two cases in the discussion of Model A. The results for all photons case are shown in Fig. 1. The left panel shows the posterior distributions and the right panel shows the fitting results where all 14 photons are denoted by gray dots and the 9 mainline photons are denoted by gray dots with orange error bars, respectively. The results suggest that the mean value of intrinsic time delay $\mu=-6.48^{+4.53}_{-4.50}$ s, and $a_{\rm LV}=2.12^{+0.57}_{-0.57}\times 10^{-18}\ \rm{GeV^{-1}}$. The corresponding $E_{\rm LV}=4.71^{+0.17}_{-0.10}\times 10^{17}$ GeV. The results for the mainline photons case are shown in Fig. 2. The results suggest that the mean value of intrinsic time delay $\mu=-10.34^{+1.07}_{-1.10}$ s, and $a_{\rm LV}=2.75^{+0.10}_{-0.11}\times 10^{-18}\ \rm{GeV^{-1}}$. The corresponding $E_{\rm LV}=3.64^{+0.01}_{-0.01}\times 10^{17}$ GeV. Results for both two cases are consistent with the results in [21, 22]. Both cases suggest that high-energy photons are emitted earlier than low-energy photons because of the negative value of $\mu$, and the LV time delay term contributes to the observed time delay.
@@ 3.3 Model C
$\begin{cases}p\left(\mu\right)\sim U\left[-30,30\right]\ {\rm s},\\ p\left(\sigma\right)\sim U\left[0,30\right]\ {\rm s},\\ p\left(a_{\rm LV}\right)\sim U\left[0,30\right]\times 10^{-18}\ \rm{GeV^{-1}},\\ p(\alpha)\sim U\left[-30,30\right]\ {\rm s\cdot GeV}.\\ \end{cases}$ (21)
As there are both $E_{s}$-dependence and $K_{1}$-dependence in the fitting procedure rather than a simple linear relation, we should fit all 14 high-energy photons in Model C. Then we obtain the results shown in Fig. 4. The left panel shows the posterior distributions and the right panel shows the fitting results. The results suggest that the mean value of intrinsic time delay $\mu=-4.98^{+4.26}_{-4.32}$ s, $a_{\rm LV}=3.38^{+1.05}_{-0.98}\times 10^{-18}\ \rm{GeV^{-1}}$, and $\alpha=-0.15^{+0.10}_{-0.11}\ {\rm s\cdot GeV^{-1}}$. The corresponding $E_{\rm LV}=2.96^{+0.12}_{-0.07}\times 10^{17}$ GeV. We obtain a physics picture similar to that of Model A. Both the common intrinsic time delay term and the source energy correlated intrinsic time delay term are negative, which suggests that the high-energy photons are emitted earlier than the low-energy photons. Meanwhile, the $E_{\rm LV}\simeq 2.96\times 10^{17}$ GeV in Model C is smaller than the results in Model A as $E_{\rm LV}\simeq 4.71\times 10^{17}$ GeV for all 14 high-energy photons case, implying a slightly stronger LV effect. It is worth mentioning that we still need more high-energy photons from GRBs to get more solid conclusions.

### v2

@@ 3.1 Model A
[…]
@@ 3.3 Model C
$\begin{cases}p\left(\mu\right)\sim U\left[-30,30\right]\ {\rm s},\\ p\left(\sigma\right)\sim U\left[0,30\right]\ {\rm s},\\ p\left(a_{\rm LV}\right)\sim U\left[0,30\right]\times 10^{-18}\ \rm{GeV^{-1}},\\ p(\alpha)\sim U\left[-30,30\right]\ {\rm s\cdot GeV^{-1}}.\\ \end{cases}$ (21)
As there are both $E_{s}$-dependence and $K_{1}$-dependence in the fitting procedure rather than a simple linear relation, we should fit all 14 high-energy photons in Model C. Then we obtain the results shown in Fig. 4. The left panel shows the posterior distributions and the right panel shows the fitting results. The results suggest that the mean value of intrinsic time delay $\mu=-4.98^{+4.26}_{-4.32}$ s, $a_{\rm LV}=3.38^{+1.05}_{-0.98}\times 10^{-18}\ \rm{GeV^{-1}}$, and $\alpha=-0.15^{+0.10}_{-0.11}\ {\rm s\cdot GeV^{-1}}$. The corresponding $E_{\rm LV}=2.96{\color[rgb]{1,0,0}{}^{+1.21}_{-0.70}}\times 10^{17}$ GeV. We obtain a physics picture similar to that of Model A. Both the common intrinsic time delay term and the source energy correlated intrinsic time delay term are negative, which suggests that the high-energy photons are emitted earlier than the low-energy photons. Meanwhile, the $E_{\rm LV}\simeq 2.96\times 10^{17}$ GeV in Model C is smaller than the results in Model A as $E_{\rm LV}\simeq 4.71\times 10^{17}$ GeV for all 14 high-energy photons case, implying a slightly stronger LV effect. It is worth mentioning that we still need more high-energy photons from GRBs to get more solid conclusions.